### v1

# Flexible Process Variant Binding in Information Systems with Software Product Line Engineering

Abstract
Different organisations often run similar digitised business processes to achieve their business goals. However, organisations often need to slightly adapt the business processes implemented in an information system in order to adopt them. Various approaches have been proposed to manage variants in process models. While these approaches mainly deal with control flow variability, in previous work we introduced an approach to manage implementation variants of digitised business processes. In this context Software Product Line (SPL) Engineering was applied to manage a set of common core artefacts including a process model from which Process-Aware Information Systems (PAIS) can be derived, which differ in the implementation of their process activities. When deriving a PAIS, implementations are selected for each process activity and then included in the PAIS at compilation time. One challenge that has not yet been solved is giving users of digitised business processes the option of selecting several features at runtime, i.e. selecting multiple activity implementations at runtime. This paper extends our previous work by not only allowing for the selection of activity implementations at compile time, but also at start time and runtime. Consequently, it becomes possible to defer the decision as to which features should be selected to start time and runtime. Furthermore, multiple implementations of a particular activity may be selected and executed concurrently. As another challenge different organisation may want to collect and base their decision on different information in a digitised business process. Consequently, the presented approach also allows customising the input and output data of activities when deriving a PAIS for a specific organisation.
Keywords:
Business Process Management Software Product Line Engineering Process Configuration Process Variability Software Reuse
1 Introduction
The right to self-administration of German municipalities leads to various variants of the same business process. For example, craftspersons may apply for a special parking permit, which allows them to park in areas in which regular citizen have to pay or in which generally parking is not permitted. Special parking permits for craftspersons are available, inter alia, in the German cities of Munich [^1] http://www.muenchen.de/dienstleistungsfinder/muenchen/1072021/, Constance [^2] https://service.konstanz.de/Leistung-Detailansicht/?id=6000803, and Stuttgart [^3] https://www.stuttgart.de/vv/leistungen/sonderparkausweise-fuer-gewerbetreibende-und-soziale-dienste.php. The business process for checking an application for the special parking permit is similar between the municipalities. However, there are slight differences among the municipalities. In some municipalities, the special parking permit is issued automatically, whereas in others a municipal employee needs to issue the parking permit manually. Other researchers have observed and investigated business process variability in municipalities as well [33], focusing on control flow variability.
Approaches have been proposed to deal with business process variability. Reference processes [52] and base processes [21], respectively, were introduced that may be configured to meet the needs of an individual organisation. However, these approaches focus on variations in the control flow of a business process rather than on the implementation level. Software Product Line (SPL) Engineering [6] addresses the challenge of developing and maintaining a set of similar software products. A Software Product Line comprises a common set of core artefacts from which features may be selected in order to build a specific software product. In literature, the process of selecting features from an SPL and building a software product is referred to as product derivation [19][28]. A Process-Aware Information System (PAIS) corresponds to a software product that executes a business process involving human actors, applications, and information sources [20]. In previous work [24], we applied and simplified the concepts of SPL Engineering to PAIS (PAIS Product Line). At build time, any implementation can be selected for an activity. However, this static approach lacks flexibility, which is of utmost importance in practice.
1.1 Problem Statement
In our previous approach, the activities of a PAIS Product Line have specified input and output data structures, i.e. it is specified what data an activity expects and what data can be collected during the execution of an activity. The combined input and output data structures of all activities are denoted as process data structure. The process data structure of the PAIS Product Line is assumed immutable, i.e. when deriving a product from a PAIS Product Line the process data structure cannot be customised. Data structures are implemented in source code (e.g. Java Classes). If the process data structure is to be customised during product derivation (e.g. a user form is adapted in order to collect different data), the corresponding source code (Java Classes) needs to be adapted, e.g. removing or adding fields in a Java Class. This poses a challenge as removing fields of a Java Class might lead to compile errors when these fields are used elsewhere.
Furthermore, the implementations for the activities need to be selected at build time. Selection at start time or runtime is not possible. In order to give the process engineer of a business process the option of selecting features, the decision as to which activity implementations are selected must be deferred to startup time or runtime. In previous work, we used the build tool Apache Maven to conditionally package the selected activity implementations. When activity implementations are to be selected at startup time or runtime a build tool cannot be used anymore. The challenge is to find and adopt a mechanism from SPL Engineering allowing for the selection of activity implementations at startup time and runtime. However, note that when it becomes possible to select and, consequently, deselect activity implementations at runtime, it needs to be specified what shall happen with the data of a running implementation that gets deselected.
Finally, so far only one implementation may be selected for an activity. In order to give the process engineer of a business process the option of selecting multiple implementations for one activity, it becomes necessary to execute the selected implementations concurrently. Concurrent running activity implementations might access the same data which can result in unintended overwriting of data.
1.2 Contribution
This work addresses the aforementioned limitations and problems of PAIS Product Lines. The main contribution of this work is threefold:
  1. 1.
    Concepts and tools from SPL Engineering are adopted and applied to the approach of PAIS Product Lines to enable customising the process data structure when deriving a PAIS product.
  2. 2.
    Techniques known from SPL Engineering are selected and combined to enable selecting activity implementations at build time, startup time, and runtime.
  3. 3.
    An approach is presented that allows selecting and running multiple activity implementations concurrently, while ensuring that no data is unintentionally overwritten.
Moreover, it is assured that the process of designing, developing, selecting activity implementations, and deriving products (known as the phases of SPL Engineering) is tool supported.
1.3 Outline
The remainder of this paper is structured as follows: Section 2 presents basic concepts of SPL Engineering, which lay the foundation of the approach and which are necessary for understanding it. The approach of PAIS Product Lines we introduced in [24] is described in Section 3. In Section 4, this approach is extended to enable more flexibility in that the data structure may be customised during PAIS Product derivation, multiple implementation for one activity may be selected, and the implementation may be bound at different times. The flexibility of the presented approach is discussed in Section 6. Section 7 discusses related work. Section 8 concludes the paper and provides and outlook on open challenges and further research topics.
2 Fundamentals
In this section, basic concepts of SPL Engineering are presented, which constitute the foundation of our approach. The concepts deal with variability among similar software products that may be derived from a set of common core artefacts.
2.1 Variability Modelling
Kang et al. [27] introduced a notation to capture the functionality of a software product from the perspective of a user. By applying this notation, it becomes possible to outline the features of a software product as well as their relationship in a tree-structured feature model. The root node represents the software product. Features of the software product are connected by edges to the root node. Features themselves may consists of other features and are inter-connected with them via edges as well. Little circles at the end of an edge refer to optional features, whereas alternative features are identified by arcs between their corresponding edges. Feature diagrams help developers to identify what has to be configured in a software product.
An example of a feature model is displayed in Figure 1a. Software Product 1 consists of three feature, i.e. Features f1, f2, and f3. While f1 is optional, f2 and f3 are mandatory. Furthermore, f2 consists of either Feature f2.1 or f2.2.
(a) Original Notation [27]
(b) Cardinalities [17]
(c) Constraint
Figure 1: Feature Model Notation
Besides optional and mandatory features, groups of alternative features may have assigned cardinalities [17][18] (or multiplicities [39], respectively). The cardinality of a feature group defines the number of features to be selected from this group. The feature model depicted in Figure 1b represents Software Product 2, which comprises four alternative features (i.e. f4, f5, f6, and f7). The cardinality ⟨2-3⟩ specifies that for Software Product 2 at least two features and at most three features have to be selected.
In [27], typical constraints between features such as requires or mutually exclusive are described in text form. It is now common practice to use a dashed arrow to indicate that a feature requires another one and a dashed double-headed arrow to illustrate that the two marked features must not be co-selected (e.g. [39], [8], and [45]). Finally, Figure 1c shows a feature model that contains a software product with three optional features (i.e. f8, f9, and f10). Due to the constraint requires, when selecting Feature f8, Feature f9 must be selected as well.
2.2 Feature Interactions
In a feature interaction [11][13], in combination multiple features might behave differently than in isolation. Feature interactions have been subject to research in many domains for a long time [11][13][12]. An example of a feature interaction is given by [11] and described in Example 2.2:
{customExample}
(Feature Interaction in Telecommunication): Consider the features call-forwarding and call-waiting on a busy phone line. If a phone consists of one of the two features, it works fine. However, the behaviour of a phone comprising both features is unclear. Either way the requirement of one of the features is not satisfied, maybe even the requirements of both.
[…]
In SPLs, the resolution of feature interaction cannot be hard-coded as during product derivation the features are selected individually obeying the constraints set out by the feature diagram [25]. However, additional code that resolves the unintended behaviour may be conditionally included when two interacting features are present in a software product as proposed for the optional-feature problem. If a feature contains code that depends on another optional feature, the second feature becomes mandatory contrary to its specification, which is known as the optional-feature problem [29]. Different approaches propose extracting the dependent code into a separate module, which is called derivative [35][36] or lifter [38]. Consequently, both features may be included independently and used in software products. However, when using both features in combination, the derivative module is included as well. A derivative module constitutes the resolution of technical dependencies and is not a feature. Consequently, it is not added in the feature model [29]. Furthermore, in order to resolve the interaction of two features they may be marked as mutually exclusive in the feature model, or precedences or priorities may be defined. For a review of resolution techniques for feature interactions we refer interested readers to [48].
2.3 FeatureHouse
FeatureHouse [4] is a framework and tool chain for composing software artefacts to software systems. It uses superimposition to merge the software artefacts, i.e. the substructures of the software artefacts are merged in order to compose the software system. The hierarchical structure of a software artefact is represented by a feature structure tree (FST). For example, a Java artefact may consist of packages, classes, and methods that correspond to FST nodes. Nodes may be non-terminals (i.e. having child nodes) or terminals. Depending on the hierarchy level at which the artefacts shall be merged, the structural elements of the latter need to be chosen as terminal nodes. If, for example, Java artefacts shall be merged at the class level, the classes may be the terminal nodes. Superimposing two FSTs entails merging their nodes beginning with the root node. FeatureHouse supports the composition of artefacts written in various languages, including Java, C#, and XML.
2.4 Feature Refinement
A feature refinement [7] corresponds to a program increment that represents a feature within related software products, e.g. a software product line. Feature refinements comprise software fragments that can be incrementally composed to build a software product whereby the software fragments may be source code and other artefacts. Jak [7] is an extension of the programming language Java that contains key words for feature refinement. A class contained in a Jak file may be refined by another Jak file using the modifier “refines”, e.g. class members and methods may be added.
The AHEAD tools presented in [7], in turn, are able to compose and translate Jak files to Java files. The AHEAD tools also support incrementally composing XML documents written in XAK [3]. XAK is a language to refine XML documents and provides the attribute ”xak:module” to mark a tag as a module (i.e. this tag may be refined) in the base XML file, which is the starting point for composing XML documents. All tags in the XML tree below the module are considered implementation and must not be refined. In separate XAK files, refinements of a module (i.e. XML increment) may be specified. During composition, the refinements are applied to their corresponding module tag in the base XML file, i.e. the tags in the XML-refinement are appended to the corresponding module tag. As XAK refinement files constitute only an increment rather than a complete XML, the files are mostly not schema-compliant.
2.5 Feature Binding
In order to use an optional feature of an SPL, it has to be selected in respective software product. This is called feature binding [16]. Features may be bound at different times. In this context, [27] distinguishes three feature binding times: compile time, start time, and runtime. Features bound at compile time are composed during source code compilation, whereas features bound at start time are selected when a software product is launched. Finally, features bound at runtime may be exchanged while the software product is running. Propositions such as [42] and [41] categorise feature binding times into static binding (i.e. feature binding before software product execution) and dynamic binding (i.e. feature binding during start or while running a software product). Static and dynamic binding each have advantages and disadvantages [10]: On the one hand, dynamic binding entails flexibility with respect to adding and removing features at runtime. On the other hand, this also implies that the software product contains all features. The size of a software product has an impact on the required resources such as memory and CPU. In contrast, static binding requires fewer resources but lacks flexibility at runtime.
Dynamic binding implies adding and removing features during runtime. Hence, it becomes necessary to specify what happens if a running feature shall be removed. [34] summarises the options:
  • The running feature is removed immediately.
  • The current state of the feature is saved before its removal.
  • The running feature is removed after its execution has completed.
  • The running feature is not removed at all.
Furthermore, the removal of a running feature might impact other features with which it shares data or system processes (i.e. background tasks of the operating system) [34]. [34] proposes grouping features into feature binding units at development time that can be bound and unbound at runtime. The feature binding units consider implications of data and system processes the features to be unbound are working on.
A variation point corresponds to the location in an information system at which variability occurs, i.e. features are bound. Methods to implement a variation point (i.e. binding a feature) are called variability mechanisms [26]. In earlier approaches, the binding time of a feature had to be selected at design time by implementing the variability with an appropriate variability mechanism [14][42][10]. For instance, the variability mechanisms discussed in Subsections 2.3 and 2.4 like FeatureHouse [4] and AHEAD [7] allow for static feature binding, solely.
The approach presented in [42] aims to support static and dynamic feature binding without deciding at design time. The code basis of the SPL is developed independently of the feature binding time by using feature refinements. If the features of the SPL shall be bound statically, existing tools are used for composing the feature refinements. In order to bind the features dynamically, the code base is transformed to use the decorator pattern. Features can be selected after compile time by using factory methods of the feature decorators. Thus, all features are included at compile time and available afterwards.
The approach allows selecting either static or dynamic feature binding for all features. In contrast, [41] proposes an approach that builds on the results of [42], but allows selecting the binding time per feature. Changing the order in which features are bound may change their behaviour. Consider Example 2.5.
{customExample}
(Feature Binding Order): Assume Feature A needs to be executed before Feature B. Then, Feature A needs to be bound after Feature B. However, if Feature A is bound statically and Feature B dynamically the order is reversed. The approach presented in [41] ensures the execution order of the features even when features have different binding times. This is achieved by generating hook methods that can be overridden.
In dynamic feature binding, mutually exclusive features are included together in the software product as they are not bound before runtime. Therefore, when binding features at runtime, given constraints (e.g. mutually exclusion and implication) need to be obeyed. Both [42] and [41] use a runtime API that enables checking against the corresponding feature model on whether or not the selected features in an SPL Product are consistent.
Edicts constitute aspects of aspect-oriented programming [14]. For a variation point there may be various edicts. If a feature shall be bound at compile time an edict containing this feature is included. In contrast, if the implementation of the variation point shall be determined at runtime, an edict containing the code of a design pattern (e.g. decorator pattern) that enables choosing a feature at runtime is included.
2.6 Phases of Software Product Lines Engineering
As opposed to single software products, when developing an SPL the requirements of various similar, yet different, software products need to be collected, managed and implemented. SPL Engineering can be divided into four phases [51][28]:
  1. 1.
    Domain analysis includes the activities of collecting the requirements and describing the domain features of the SPL, i.e. collecting the requirements for software products that can be derived from the SPL.
  2. 2.
    Domain implementation deals with implementing the features described during domain analysis in a way that allows composing them individually for each derived software product.
  3. 3.
    Requirements analysis deals with selecting the features from the domain analysis for a specific software product derived from the SPL. The selected features form a configuration.
  4. 4.
    Software generation deals with building a software product by composing those features specified in the configuration from the requirements analysis.
2.7 FeatureIDE
FeatureIDE [51] is an Eclipse-based integrated development environment (IDE) for SPLs. It covers the four phases of Software Product Line Engineering set out in Section 2.6). To support domain analysis, FeatureIDE provides a graphical editor to model features and their dependencies in a feature model. Feature models contain constraints, are stored as XML files, and can be imported to or exported from FeatureIDE. Furthermore, FeatureIDE provides the user with convenient refactoring tools and inconsistency detection.
FeatureIDE assists the developer with a configuration editor to create a configuration for a specific software product during requirement analysis. The developer may mark those features from the feature diagram to be contained in the software product. The configuration editor prevents configurations not obeying the constraints of the feature diagram. Finally, the configuration is stored in a configuration file.
FeatureIDE supports various frameworks for implementing features during domain implementation. This includes FeatureHouse and the AHEAD tools.
For software generation, FeatureIDE takes a configuration file as input and composes the software artefacts corresponding to the selected features. FeatureIDE places the composed software artefacts in a specified output directory. The artefacts in the output directory are then natively compiled as artefacts of the language the framework for the domain implementation is based on.
3 Previous Work
In our previous work, we introduced the approach of Process-Aware Information System Product Line (PAIS Product Line) [24]. This approach applies the concepts of SPL Engineering to the development and maintenance of similar PAISs in order to reduce development efforts and costs by enabling the reuse of common core artefacts.
3.1 Variation Points
A PAIS Product Line constitutes a common set of core artefacts including a core process model from which PAIS products may be derived. Various activities of the core process model may be declared as variation points, i.e. different implementations may be selected for these activities during PAIS Product derivation. In BPMN 2.0444https://www.omg.org/spec/BPMN/2.0/PDF, an activity corresponds to a step on which work is performed, which is either atomic or compound. While call activities (call subprocess, which itself comprise activities) constitute compound activities, tasks are atomic activities that cannot be further broken down. For the sake of this approach, the BPMN 2.0 send task, script task, service task, and business rule task are regarded as automated tasks as they all represent an automatic step in a business process without user interaction.
Neither manual tasks nor the receive tasks are considered by this approach as manual tasks do not occur in digitized business processes and receive tasks may be substituted by receive events. A BPMN 2.0 user task corresponds to a step in a business process where the user interacts with the PAIS. Consequently, the approach considers the following three activity types: call activity, automated task, and user task.
Figure 2 shows an example business process comprising two activities, whose activity types are not specified, i.e. the activities constitute the variations points. During PAIS product derivation, an implementation will be selected for these activities and hence the activity types be determined. Depending on the selected implementation, the activities might be automated activities (i.e. executing business logic), user tasks (i.e. a user form) or call activities (i.e. calling a subprocess). An activity might also remain without a type, i.e. no implementation is selected. Consequently, the activity is neither a call activity nor an automated task nor a user task. An untyped activity has no function in a PAIS. A process instance that passes an untyped activity does nothing, i.e. neither a user form is invoked nor automated logic is executed.
Figure 2: Example Business Process [24]
3.2 Variability Modelling
The approach transfers the concept of features from SPL Engineering to process activities. This enables us to create process features models, which depict what implementations may be selected for which activity, and which activities are optional and which mandatory. The process feature model depicted in Figure 3 reflects the process model from Figure 2. In the process feature model, for each of the two activities one out of two implementations may be selected. Furthermore, Activity 2 is optional, i.e. an implementation may be selected, but this is not mandatory.
A process feature model consists of three layers, i.e.. the process level, the feature level, and the implementation level. The process level represents the entire business process. The feature level includes all activities of the business process and specifies whether they are mandatory (i.e. an implementation has to be selected) or optional (i.e. an implementation may be selected but this is not mandatory). The implementation level lists the available implementations of the activities. If an activity corresponds to a subprocess, another process feature model with the same structure (3 levels) has to be constructed for this subprocess.
Figure 3: Process Feature Model [24]
3.3 Feature Interaction
Each activity of a business process has input and output data whose data structure is specified at design time. The term process data structure is used to refer to the combined input and output data of all activities of a business process. Thus, the input/output data structure of an activity determines that part of the process data structure to which the activity has read/write access. Two activities with the same output data structure have the same write access. Hence, an activity that has the same output data structure as its preceding activity might overwrite the data of its predecessor, i.e. the data of the predecessor is lost. During requirements engineering, the order of the activities as well as their output data structure (i.e. write access) is specified. Therefore, the behaviour of the activities can be foreseen and data will not inadvertently be overwritten. Aside from passing on and accepting data, there is no interaction between the activities as they are executed in a predefined order and not in parallel. Consequently, there are no unintended feature interactions in PAIS Product Lines.
4 Flexibility for PAIS Product Lines
The approach of PAIS Product Lines currently lacks flexibility. Only one implementation can be chosen for an activity during PAIS Product derivation. Furthermore, the process data structure cannot be adapted to the individual need of an organisation. All derived PAIS Products share the same process data structure. Finally, the approach of PAIS Product Lines currently supports feature binding at compile time, solely.
This section shows how to eliminate the limitations to flexibility by enhancing the approach of PAIS Product Lines. First, the requirements of the enhanced approach are elicited. Then, the assumptions for the enhanced approach are stated. Finally, it is shown how the aforementioned limitations concerning flexibility can be eliminated.
4.1 Requirements
In [24], we used the business process special parking permit for craftspersons in a simplified version as an example, which we deduced from a cooperation with German municipalities. In this work we extend the example towards a more complete version of the business process. Example 4.1 describes the business process.
{customExample}
(Special Parking Permit Process): Various German municipalities offer a special parking permit for craftsperson who may park their cars in zones where regular citizens need to pay or where parking is prohibited. The business process of checking an application for this parking permit is the same among all considered municipalities and is displayed in Figure 4. First, a craftsperson applies for the parking permit. Then, the application is checked. If the application is justified, the parking permit is issued. If the application is not justified, the applicant is notified of the rejection.
While the described control flow is common across all municipalities, the implementation of the business process varies. In some municipalities, the application check is carried out by a municipal clerk whereas in other municipalities the application is checked automatically by comparing the data of the application with the data of the craftsperson who is officially registered with the authorities. Besides, when the application is rejected the applicant may be notified via e-mail, or sms, or by a municipal clerk depending on the municipality and the choice of the applicant. The applicant might choose the way she will be notified from a list of available notification means.
Furthermore, the process data structure also varies as the municipalities require the applicant to fill in different data. In some municipalities, the parking permit is valid for every car whereas in others the applicant needs to provide the number plate for which the parking permit shall be valid in order to prevent misuse of the parking permit for the craftsperson’s private car. In addition, to be able to automatically check the application, the craftsperson needs to provide the commercial register number under which she is registered with the authorities.
Figure 4: Business Process Special Parking Permit for Craftspersons
In a nutshell, PAIS Product Lines need to comprise common core artefacts including a process model from which similar PAIS Products can be derived in order to reduce the development efforts in comparison to developing the similar PAIS Products separately. One or more implementations shall be selectable for an activity that constitutes a variation point. The implementations shall be selectable at compile time, start time, and runtime and the process data structure shall be customisable for each PAIS Product as well. The business process of checking an application for the special parking permit and the resulting requirements were studied in a cooperation with German municipalities. While in previous work we used a simplified version of the business process as an example, in this work we use a more comprehensive version and deduce the requirements for our enhanced approach accordingly. The enhanced approach of PAIS Product Lines shall meet the following requirements:
  1. R1:
    PAIS Products need to be derivable from a PAIS Product Line which denotes a set of common core artefacts including a core process model in order to avoid the development of redundant software artefacts for similar PAISs.
  2. R2:
    Multiple implementations for one activity shall be selectable.
  3. R3:
    During PAIS Product derivation the process data structure shall be adaptable allowing different organisation to use different process data structure in their derived PAIS Product.
  4. R4:
    It needs to be formally specified what implementations for which activities may be selectable and what constraints need to be met during this selection.
  5. R5:
    Combining certain implementations of different activities might produce unintended behaviour (feature interaction). Methods need to be established in order to detect and prevent unintended feature interactions.
  6. R6:
    To enable selecting activity implementations during start and runtime based on input of users dynamic feature binding becomes necessary. When using dynamic feature binding, the challenges described in Section 2.5 need to be tackled:
    1. (a)
      When an activity implementation is dynamically unbound during runtime (i.e. the implementation is deselected) the data it was processing might get lost. It needs to be specified whether an activity implementation may be unbound and - if yes - how, e.g. saving the current state.
    2. (b)
      When unbinding an activity implementation during runtime, shared data and system processes might be affected. Methods need to be established ensuring that shared data and system processes will not be corrupted.
    3. (c)
      When using static and dynamic feature binding, activity implementations might be bound in a different order than intended. It must therefore be assured that different binding orders must not lead to unintended behaviours.
4.2 Assumptions
For the approach to work, the following assumptions hold:
  • The control flow of the core process model included in a PAIS Product Line is specified and immutable, i.e. the process model does not change during PAIS Product derivation. Hence, all derived PAIS Products will have the same process model and thus an identical control flow. While this work focuses on implementations variability, other approaches deal with control flow variability [52, 21, 23, 22].
  • The concepts will be applied to activity-centric business processes. Object- and data-centric approaches to business process management like those discussed in [30] and [49] are excluded in this approach.
4.3 Multiple Implementations of an Activity
In our previous work, when deriving a PAIS Product, for each activity constituting a variation point only one implementation may be selected, i.e. the implementation may be selected if this activity is either a manual task consisting of a user form or an automated task executing business logic. In Figure 5 an extract of the business process special parking permit is depicted. Activity Check application has no type and hence corresponds to a variation point. Two implementations are available, i.e. Automatic Check and Manual Check. In our previous work, only one of the two implementations may be selected. Furthermore, the activities of a business process are executed in a pre-specified order, e.g. activity Check application is executed after activity Apply for special parking permit. The execution order as well as the read and write access of each activity to attributes of the process data structure is determined during requirements engineering. In Figure 5, it can be seen that activity Apply for special parking permit has write access to the Application data structure, whereas activity Check application has read access to Application and write access to the Decision data structure. Consequently, the activities only influence each other with the data they read and write. As the read and write access is carefully specified there are no unintended feature interactions.
Figure 5: Extract of Business Process Special Parking Permit for Craftspersons with Implementation Options
When allowing for multiple implementations of an activity, no execution order of these implementations is specified. If, for example, in Figure 5, for activity Check application both implementations Automatic Check and Manual Check are selected, both implementations run in parallel when the process instance passes this activity. As the implementations of an activity may be executed in isolation there is no unintended technical feature interaction. However, implementations of one activity may influence each other via the data they write. Due to non-determinism it might not been foreseeable which implementation finishes last and thereby overwrites the data of its predecessors, i.e. the results are unpredictable. Consider Example 4.3.
{customExample}
(Feature Interactions in PAIS Product Lines): In Figure 5, activity Check application writes the decision on whether an application is justified to the process data structure. Consequently, both implementations may write the decision as well, i.e. the application is checked both automatically and by a municipal clerk. When selecting both implementations (i.e. manual and automatic) for activity Check application, the implementation that finishes last determines whether the application is justified. For example, the result of the automatic application check is overwritten by the decision of the municipal clerk if he/she takes longer than the automatic application check. This leads to unpredictable results depending on which implementation writes the data last. This can be seen as unintended feature interaction.
To tackle this challenge, activities reading and those writing need to be distinguished. Activities that have read-only access to data do not pose a challenge in terms of feature interaction. In Example 4.1, when selecting both sms and e-mail for notifying the applicant it does not matter what implementation finishes first as the implementations do not write data and hence do not influence each other. In order to prevent feature interactions in activities that write data, prevention techniques from SPL Engineering might be taken into consideration.
In Section 2.2, mutual exclusion of the interacting features, precedences and priorities are discussed as solutions to prevent feature interactions. However, due to Requirement R2 it becomes necessary to be able to co-select multiple implementations for one activity, mutual exclusion cannot be used to prevent feature interaction. Specifying precedences and priorities means that the implementation with the highest priority will always overwrite the data of the other implementations. Consequently, there is no sense in including multiple implementations when the one with the highest priority always overwrites the data of the others.
Another option would be to not grant the different implementations write access to the same data. This does not work either as the implementations implement the same business requirement in different ways and therefore need to write the same data. For example, the application check implementations all need to write the data on whether or not the application is justified. Note that the concept of derivative modules (cf. Section 2.2) may be used in this context as well.
Additional code can be included when two implementations write to the same data structure. For example, additional code could determine that the application is not justified as soon as at least one of the two implementations decides that the application is not justified. This additional code aggregates the data from the different implementations. We refer to this additional code as aggregation code. Another aggregation code might constitute a majority vote. If there are multiple application check implementations, the application is rejected when a majority of the implementations consider the application being unjustified. As opposed to derivative modules, aggregation code needs to be placed in the process feature model as aggregation code does not solely resolve technical dependencies. Aggregation code introduces resolution on a business level. There might be multiple aggregation codes from which one may be selected during PAIS product derivation. This choice needs to be indicated in the process feature model. In the process feature model all implementations of one activity are grouped as well as all available aggregation codes for that activity. A dashed arrow introduces the constraint that one aggregation code needs to be selected when more than one implementation of an activity is selected. Consider the following example.
{customExample}
(Aggregation Code): Figure 7 shows the process feature model for Example 4.1. For activity Check application two implementations may be selected, a form for a manual check and/or an automatic check. In addition, two aggregation code implementations are available. The feature model notation described in Section 2.1 was extended in Figure 7. The two implementations for activity Check application as well as the two aggregation code implementations are each grouped using a dashed box.
A dashed arrow connects the implementation group I and the aggregation code group A of the same activity. Its label states
#$\#I>1?requires\#A=1$
which reads as follows: If the number of selected implementations is greater than one, one aggregation code implementation needs to be selected, i.e. if two items from Group I are selected, one item from Group A needs to be co-selected as well. The dashed arrow, which represents a conditional requires, was taken from literature (cf. Section 2.1) and supplemented with a label.
4.4 Varying Data Structures
In our previous work, during PAIS Product derivation the process data structure could not be customised to the individual needs of an organisation. All derived PAIS Products shared the same process data structure. However, Example 4.1 describes a scenario in which it becomes necessary to adapt the process data structure during PAIS Product derivation.
{customExample}
(Varying Process Data Structure in a PAIS Product Line): When applying for the special parking permit, some municipalities require the applicant to fill in the number plate whereas the parking permit of other municipalities is valid for all vehicles it is placed in. Furthermore, to automatically check the application, the application requires the commercial register number. Consequently, the process data structure varies between the municipalities.
As can be seen in Example 4.4, different organisations need to customise the process data structure when deriving a PAIS Product from a PAIS Product Line. Consequently, there are optional parts of the process data structure. Activities read and write data. The optional parts of the process data structure are accessed at least by some activity implementation or otherwise the selection of the former was useless as they are never used. Hence, there are dependencies between activity implementations and the data structure. Including the process data structure with the optional parts in the process feature model allows pointing out the dependencies between the activity implementations and the optional parts of the process data structure. To point out the dependencies dashed arrows like the ones used in literature can be employed. During PAIS Product derivation, when the activity implementations are selected, the corresponding optional parts of the process data structure need to be included in the PAIS Product in order to enable the compilation of the latter. From a business requirements perspective, each included optional part of the data structure needs to be accessed at least once for writing and once for reading, in exactly this order. Otherwise, if one part of the data structure is accessed for reading before writing, the data structure will be empty. Furthermore, if one part of the data structure is accessed for writing, solely, there is no point in collecting the data at all as the data is never being processed or used.
In the following, the process of checking an application for the special parking permit (cf. Examples 4.1 and 4.4) is used to explain the concept of varying data structures. Figure 6 shows the process data structure associated with Example 4.4 as UML class diagram. The application corresponds to a composition of the data of the applicant, the data of the company, and the car information, whereas the latter and the field commercial register number in the data of the company are optional (indicated with question marks). The optional parts of the data structures are only included if necessary.
Figure 6: Data Structure of Example 4.4
Each activity implementation requires access to a specific part of the data structure, e.g. the automatic application check requires the commercial register number field in the data of the company. In addition to the process feature model, bottom of Figure 7 contains a feature model representing the data structure from Figure 6. Those activity implementations that need a specific optional part of the data structure have dashed arrows to the required data structure part with a label indicating whether the implementation needs read or write access. For activity Apply for parking permit multiple alternative implementations exist. There are three forms: a simple form, a form in which the applicant needs to fill in the commercial register number, and a complex form that contains input fields for setting the commercial register number and the number plate. All forms need write access to the corresponding data structure parts. The automatic application check implementation needs read access to the commercial register number to be able to compare the application data with the data from the commercial register. For activity Issue parking permit, there are two alternative forms. One does not require optional parts of the data structure, whereas the other form requires read access to the number plate field as the number plate needs to be present on the issued parking permit.
Figure 7: Process Feature Model
4.5 Static and Dynamic Feature Binding
The activity implementations are independent from each other and have no technical inter-dependencies. The single point of interaction among the activity implementations is the write and read access to the process data structure. Consequently, the order in which the activity implementations are bound does not influence their behaviour.
If an activity implementation runs during process execution, as it was bound during process start and selected by a user, the implementation will not be unbound during execution. Once selected for a process instance, an activity implementation will always finish it execution. Hence, no data will be lost.
Shared data and system processes do not pose a problem as activity implementations are independent and have no technical dependencies. Furthermore, activity implementations may not be unbound during process execution once they are selected.
5 Proof of Concept
To validate the presented approach, Example 4.1 is implemented as a proof of concept. The Camunda Platform is used as Workflow Management System. [^5] https://docs.camunda.org/manual/7.19/ First, the properties are stated that need to be provided by the proof of concepts. Then, from a technical perspective the proof of concept shall show how to customize the process data structure during PAIS Product derivation, how to bind multiple implementations for one activity, and how to bind the implementations at different times. Finally, it is portrayed how the development and maintenance may be supported by a tool.
5.1 Properties
In line with Requirements R1-R6 (cf. Section 4.1), the proof of concept needs to provide the following properties. In addition, the development phases of a PAIS Product Line shall be tool-supported similar to SPL Engineering, which results in P7.
  1. P1:
    PAIS Products are derivable from a PAIS Product Line. All derivable PAIS Products include the mandatory activity implementations, but may differ in the selection of alternative activity implementations. The proof of concept has to show how PAIS Products can be derived from a PAIS Product Line from a technical perspective, i.e. what variability mechanism is used.
  2. P2:
    Multiple implementation of an activity might be selected and executed. A suitable mechanism needs to be chosen that allows to execute multiple implementations for one activity.
  3. P3:
    A variability mechanisms needs to be used to compose the process data structure during PAIS Product derivation to meet the organisations’ individuals needs.
  4. P4:
    A process feature model outlines all activities (incl. implementations). Furthermore, it illustrates whether an activity is mandatory or optional.
  5. P5:
    When using multiple implementations for one activity, their output data must be collected and aggregated in order to prevent the overwriting of data. This aggregation code needs to be conditionally included when more than one implementation of an activity is selected.
  6. P6:
    In order to support both static and dynamic feature binding adequate variability mechanisms must be chosen.
  7. P7:
    The phases domain analysis, domain implementation, requirements analysis, and product derivation shall be tool supported.
5.2 Data Structure Composition
In our previous work, the build tool Apache Maven was used to compose the artefacts during PAIS Product derivation. The data structure was the same among all PAIS Products and the activity implementations were designed as independent logical units and were compiled JARs (Java applications) that were conditionally included when the implementation was selected based on Apache Maven profiles. That is, Apache Maven was used to compose fully compiled artefacts. Note that this worked, as these logical units were self-contained and could be put together like building blocks. In contrast, when the process data structure needs to be customizable (cf. Example 4.4), there are changes on the class level, i.e. before its compilation the process data structure needs to be customised. Therefore, instead of composing JARs, Java classes need to be composed. To compose Java classes, both FeatureHouse [4] and the AHEAD tools [7] may be used. As FeatureHouse does not require any additional key words like refines, FeatureHouse is used to compose the Java classes of the process data structure in the proof of concept.
In Figure 6, class Company has the optional field commercialRegisterNo and the parent class Application has an optional association with the class CarInformation.
Listing 1: Base Class “Company”
public class Company {
private String name;
private String street;
private String city;
private String zipCode;
Listing 2: Classs “Company” Extension
public class Company {
private String commercialRegisterNo;
Listing 3: Composed Classs “Company”
public class Company {
private String name;
private String street;
private String city;
private String zipCode;
private String commercialRegisterNo;
Listing 1 shows the base Java class Company with its attributes. The extension of class Company with the attribute commercialRegisterNo is depicted in Listing 2. For the sake of brevity, the getter and setter methods of the classes in the listings are omitted. The class composed by FeatureHouse from the base class and its extension is shown in Listing 3. The base class, its extension, and the composed class are all pure Java classes.
5.3 Multiple Implementations of an Activity
Specific activities shall constitute a variation point, i.e. one or more implementations shall be selectable for these activities. An implementation may be a service task, a user task, or a call activity, which calls a subprocess. During PAIS Product derivation, the core process model needs to be composed with the implementations. In this context, the implementations may be considered as refinements of the base process model. The language Xak [3] can be used to compose the process models. Note that the latter are modelled in terms of BPMN 2.0, which has an XML representation.
However, when using Xak, the refinements (i.e. the implementations) are not schema-compliant, i.e. the files in which an activity is refined (which are process models as well) cannot be edited in or displayed by BPMN 2.0 modelling tools. As one advantage, BPMN 2.0 has both an XML and a graphical representation of the process model. The graphical representation serves as a means of communicating process models between domain and IT experts. If the process models serving as implementations are not schema-compliant, the advantage of BPMN 2.0, i.e. the graphical representation, does not come into effect. Furthermore, when selecting lots of implementations these implementations need to be included in the core process model and the elements need to be rearranged in order to make room for the additional elements of the implementations. Although works such as [46] provide means for automatically layouting process models, with increasing number of included implementations the process model gets unclear and confusing. Therefore, Xak is not used to compose the base process model and its implementations. Instead, multi-instance call activities are used, which call the implementations as subprocesses. The implementation (i.e. automated task or user task) are embedded in a process model, which is then called by the variation point activity. Figure 8 shows a simple process with a variation point activity which corresponds to a multi-instance call activity that calls two implementations. The implementations, in turn, are simple processes consisting of an activity, which is either an automated task or a user task. As the variation point is a multi-instance activity it may invoke both implementations in parallel allowing for the execution of multiple implementations at the same time.
Figure 8: Call Activity Calling Multiple Implementations
Figure 9 shows the process model of Example 4.1 for the proof of concept. Activities Check application and Notify craftsperson are variation points, for which one or more implementations may be selected. Therefore these activities are multi-instance call activities, i.e. they call the corresponding implementations as subprocesses. A detailed description on how the implementations are bound and called by the core process model is provided in Section 5.4.
Figure 9: Business Process Special Parking Permit for Craftspersons
5.4 Static and Dynamic Feature Binding
To be able to bind activity implementations statically as well as dynamically, approaches like the ones presented in Section 2.5 might be used. Activity implementation code might be composed during compilation when the implementations are statically bound or language constructs (e.g. design patterns like decorators) are generated that allow choosing an implementation at runtime. However, activities consist of graphical/XML representation and business logic (for service tasks) or HTML code (for user tasks). Therefore, design patterns like the decorator pattern cannot be applied to the language mixture of a PAIS Product Line. Instead, a plugin approach is used. The activity implementations (e.g. automated task or user task) are embedded in a process model consisting of a start event, an end event and the actual activity implementation. We refer to those process models embedding an activity implementation as implementation process models. Figure 8 shows a core process model (at the top) with a variation point activity and two implementation process models (at the bottom) that can be invoked by the variation point activity.
The plugin mechanism allowing for static and dynamic feature binding in PAIS Product Lines is illustrated in Figure 10. During compilation, implementation process models may be included or excluded (static binding, cf. Figure 10 (1)). Each implementation process model has a unique identifier (ID), which is used to unambiguously identify the former. When an implementation process model is included during compilation, it is registered as a plugin using its ID with the corresponding variation point activity (cf. Figure 10 (2)). The excluded implementations cannot be registered as plugins and are not available at runtime. Implementation Automatic2 in Figure 10 is not included during compilation and consequently not registered as plugin. At start time, developers and, at runtime, users may deselect one or more implementations that have been registered for a variation point activity by setting parameters. Figure 10 (3) shows a user form that can be used to start the process by selecting/excluding specific plugins (i.e. implementations). Before a variation point activity calls the registered implementation process models it is checked whether one of the latter is excluded via a parameter (cf. Figure 10 (4)). Only implementation process models are invoked that are not excluded via parameters (dynamic binding).
Figure 10: Process of Static and Dynamic Feature Binding in PAIS Product Lines
When selecting multiple implementations of a variation point activity that writes data, aggregation code needs to be included as well in order to prevent unintended overwriting of the same data. As activity implementations are realised as implementation process models, aggregation code needs to map the data returned from the subprocesses (i.e. implementation process models) to the parent process model (i.e. the core process model) by aggregating and consolidating the data from the different implementations.
From a technical point of view, for static binding, FeatureHouse is used for composing the activity implementations (like for process data structure composition). The core process is bundled with the implementation process models and the implementing logic, namely HTML code for user tasks and Java code for automated tasks. After bundling the files in one folder, the PAIS Product can be compiled. At both start time and runtime, parameters may prevent certain implementations from being selected and executed. In conditional statements in software programs, frequently parameters are used to decide whether or not a specific feature is executed based on the value of the parameter. By providing values, parameters can be exploited as variability mechanism for feature binding at runtime. Note that the use of parameters as variability mechanisms are well known in literature [2][26].
Examples 10 and 5.4 are used to demonstrate how static and dynamic binding of activity implementations works by registering them as plugins. {customExample} (Feature Binding Time in PAIS Product Lines): Activity Notify craftsperson in Figure 9 of business process special parking permit for craftsperson is used to notify the applicant when his/her application was rejected. This activity is a variation point for which multiple implementations are available. The applicant may be notified via sms, e-mail, or manually by a clerk. At compile time and start time, implementations may be selected by a developer. At runtime, the applicant can decide how he/she wants to be notified from the available notification means that have been included by a developer.
Figure 11 shows the configuration of activity Notify craftsperson for the Camunda Platform. As aforementioned, activity implementations correspond to an implementation process model that embeds the implementing functionality (i.e. automated task or user task) and the corresponding code base, i.e. Java code to implement an automated task and HTML code for user tasks. The Spring Service NotificationPluginProvider provides the IDs of the implementation process models of all registered activity implementations. Then, the Camunda Platform process engine iterates over the list of IDs using the iteration variable process and calls the implementations process models.
Figure 11: Camunda Configuration Activity Notify craftsperson
The Spring Service NotificationPluginProvider is depicted in Listing 4. It contains a field notificationPlugins of type NotificationPluginRegistration, which contains all registered implementations for notification activity. When applying for the parking permit, the applicant may select one or more notification means from all registered ones. The selection of the applicant is persisted in the process data structure. When calling the method getProcessIds (Listing 4), all registered notification plugins are retrieved, iterated over, and checked whether the applicant has selected them. The method returns the IDs of the implementation process models that are registered and the user has selected.
Listing 4: Notification Plugin Provider
1@Service
2public class NotificationPluginProvider {
3
4 @Autowired
5 private NotificationPluginRegistration notificationPlugins;
6
7 public List<String> getProcessIds(DelegateExecution execution) {
8 List<String> notificationWays = (List<String>) execution
9 .getVariable(”notificationWays”);
10 List<String> processIds = new ArrayList<String>();
11
12 for (Entry<String, String> plugin : notificationPlugins
13 .getPluginProcessIdsByPluginId().entrySet()) {
14 String pluginId = plugin.getKey();
15
16 if (notificationWays.contains(pluginId)) {
17 String processId = plugin.getValue();
18 processIds.add(processId);
19 }
20 }
21 return processIds;
22 }
23}
The Spring Service NotificationPluginRegistration (Listing 5) has the field plugins of type list, which picks up and injects all Spring Services that implement the interface NotificationPlugin illustrated in Listing 6.
Listing 5: Notification Plugin Registration
1@Service
2public class NotificationPluginRegistration {
3
4 @Value(”\${excludeNotification:}”)
5 private List<String> excludedPlugins = new ArrayList<String>();
6
7 @Autowired
8 private List<NotificationPlugin> plugins;
9
10 public Map<String, String> getPluginLabelsById() {
11
12 Map<String, String> pluginLabelsById = new HashMap<String, String>();
13 for (NotificationPlugin plugin : plugins) {
14 if (!excludedPlugins.contains(plugin.getId())) {
15 pluginLabelsById.put(plugin.getId(), plugin.getLabel());
16 }
17 }
18 return pluginLabelsById;
19 }
20
21 public Map<String, String> getPluginProcessIdsByPluginId() {
22
23 Map<String, String> pluginProcessIdsByPluginId = new HashMap<String, String>();
24 for (NotificationPlugin plugin : plugins) {
25 if (!excludedPlugins.contains(plugin.getId())) {
26 pluginProcessIdsByPluginId.put(plugin.getId(), plugin.getProcessId());
27 }
28 }
29
30 return pluginProcessIdsByPluginId;
31 }
32}
Listing 6: Notification Plugin Interface
1public interface NotificationPlugin {
2
3 public String getId();
4 public String getProcessId();
5 public String getLabel();
6}
Every implementation of the notification activity needs to implement this interface in order to get registered as plugin. The interface contains methods to get the ID of the plugin, the ID of the associated implementation process model, and the label of the plugin which is displayed in the application form (cf. Figure 12). Furthermore, the Spring Service NotificationPluginRegistration has the field excludedPlugins, which picks parameters up from command line during application start. For example, the following command starts the Spring Boot application using Apache Maven and specifies that the e-mail notification plugin shall be excluded and thereby not selectable in the application form.
mvn spring-boot:run -Dspring-boot.run.arguments=
excludeNotification=notification.mail
The Spring Service NotificationPluginRegistration comprises two methods. The first method iterates over the registered plugins, eliminates the excluded plugins, and collects and returns the labels of the remaining plugins. This method is called via the Graphical User Interface (application form, cf. Figure 12), which offers the applicant a selection of the available notification means. The second method also iterates over the registered plugins, eliminates the excluded plugins, and collects and returns the IDs of the implementation process models of the remaining plugins. The second method is called from the Spring Service NotificationPluginProvider.
Figure 12 shows an extract of the application form for the special parking permit for craftspersons. It can be observed that the applicant can choose the notification means. E-mail, sms, and notification by a clerk are available for selection in this extract, i.e. all plugins are included during compilation, are registered, and were not deselected during startup time. This Graphical User Interfaces calls the Spring Service NotificationPluginRegistration, gets the labels of the registered plugins and displays them. When the applicant submits the application form, the IDs of the selected notification means are persisted in the process data structure such that only the selected notification means are used when notifying the applicant.
Figure 12: Extract of Application Form for Special Parking Permit
{customExample}
(Aggregation Codes for Special Parking Permit Process): Activity Check application data in Figure 9, which depicts business process special parking permit for craftsperson, is used to check whether the parking permit shall be granted. This activity constitutes a variation point for which there are two available implementations. The application can be checked automatically by comparing the data from the application with the data from the commercial register. Besides, the application can be checked manually by a clerk. Both implementations write the result of the check to the process data structure. Consequently, if both implementations are selected the result data needs to be aggregated, i.e. aggregation code needs to be included that rejects the application if at least one implementation rejects the application.
When using subprocesses in Camunda, variable mappers are implemented that specify what data and how the data from the subprocess is transferred to the parent process. If only one implementation for activity Check application data is selected during PAIS Product derivation, a simple variable mapper is included that passes field justified from the subprocess to the parent process. If both activity implementations are included, a variable mapper is included that collects the fields justified from all activity implementations and evaluates them. As soon as one activity implementation finds the application unjustified the variable mapper will pass unjustified to the parent process. A variable mapper is a Java class which is referenced in the process model by its Spring bean name and may be included using FeatureHouse. In Figure 13, the configuration for activity Check application data is shown. The IDs of the implementation process models that have been included during compile time are provided by a Spring bean and then used to call the implementation process models. Furthermore, in section Delegate Variable Mapping of the configuration the Spring bean applicationCheckVariableMapper is referenced which acts as aggregation code.
Figure 13: Camunda Configuration Activity Check application data
In summary, for every activity that constitutes a variation point there is a plugin mechanism. Every implementation needs to implement the interface of the activity it wants to register with. At compile time (i.e. static binding) activity implementations may be included or excluded using FeatureHouse. The activity implementations consists of the implementation process model and the associated code base (Java code for automated tasks and html code for user tasks). At start time, command line parameters, and, at runtime, user input may be used to deselect some activity implementations (dynamic binding).
5.5 Tool Support
To support the phases of SPL Engineering with a tool, FeatureIDE is used in the proof of concept. A feature model (cf. Figure 14) visualises the variability of the PAIS Product Line during domain analysis. Note that the feature model in Figure 14 is structured from left to right such that it fits on the page. FeatureIDE captures the constraints in text form below the feature model as opposed to arrows used in other approaches.
Figure 14: Feature Model in FeatureIDE
During requirements analysis, FeatureIDE supports the developer by providing a configuration editor (cf. Figure 15). In this configuration editor the developer can select the desired activity implementations that shall be included in the PAIS Product. For implementing the variability of the PAIS Product Line, FeatureHouse was chosen, which is supported by FeatureIDE. When deriving a PAIS Product, FeatureIDE uses the the selection from the configuration editor as input (Figure 15) to compose the variable activity implementations.
Figure 15: Feature Configuration in FeatureIDE
6 Discussion
The presented approach has shown how to derive products from PAIS Product Lines with varying process data structures, which use multiple implementations for one activity, and that can bind features at compile time, start time, and runtime. In the following, the approach is assessed in respect to whether it meets the requirements set out in Section 4.1. Furthermore, the proof of concept is tested to see whether it provides the listed properties. Finally, limitations are disclosed.
6.1 Requirements
A PAIS Product Line corresponds to a set of common core artefacts from which PAIS Products may be derived. Hence, Requirement R1 is satified. Multiple implementations can be selected for one activity. When the activity may write to the process data structure, aggregation code needs to be included in order to prevent the implementations from overwriting each others data. Consequently, Requirement R2 is met. Activity implementations may require adaptions to the process data structure. When selecting a specific activity implementation the required part of the data structure needs to be selected as well. Therefore, Requirement R3 is satisfied. Requirement R4 is met as a process feature model represents the variability of a PAIS Product Line with the corresponding constraints. Unintended feature interaction can occur only when the activity implementations write to the process data structure. As aggregation code prevents implementations from overwriting each others data Requirement R5 is satisfied. At runtime, activity implementations are selected per process instance. Once they are selected, they cannot be unselected anymore. Furthermore, activity implementations can be run independently and have no technical dependency among each other. Consequently, activity implementations do not share system processes or data that might be corrupt. In addition, data does not get lost due to terminated activity implementations. Hence, Requirement R6 is satisfied as well.
6.2 Properties
The PAIS Product Line of the proof of concept uses FeatureHouse to include optional activity implementations. When including an activity implementation it registers as plugin with the corresponding activity. Consequently, the proof of concept provides Properties P1 and P2. FeatureHouse is used to compose the process data structure as well. Hence, Property P3 is provided. As a process feature model outlines the variability of the PAIS Product Line Property, P4 is provided. Aggregation code is implemented as a Camunda variable mapper class, which is conditionally included using FeatureHouse. Therefore, Property P5 is provided. Static and dynamic feature binding is supported by using a mix of FeatureHouse and parameters. Thus, Property P6 is provided. Since FeatureIDE is used as tool support in the proof of concept Property P7 is provided.
6.3 Limitations
Using FeatureHouse, each feature corresponds to a folder, which contains the code of the feature. As opposed to our previous work, the code of each feature (feature folder) cannot be compiled on its own. When deriving a product, the code of the features is composed and after the composition the code is compiled. Consequently, it is hard to develop the features independently. Although each feature folder might constitute a repository under version control, it is difficult to reference a compatible version of this feature as it is no compiled artefact with a version.
Furthermore, FeatureIDE currently lacks some convenience features. Code completion is only available for selected features and there is no package view of the feature folders. This is especially cumbersome for long package names as they are displayed as hierarchical folders instead of grouped packages.
7 Related Work
In the following, related work is assessed.
7.1 Single- and Multi-Model Approach
One way to deal with multiple variants of a business process is to include all variants in one process model by linking them with gateways [23]. However, process models containing multiple variants of the same business process may become very large and are thereby complex, unclear, and hard to maintain.
Besides combining variants into one process model, multiple process models for the various variants may be maintained [23]. Thus, similarities do not have to be investigated and the process variants can be independently modelled and evolve. A challenge is to propagate a change to all process models, which is tedious and error-prone.
The aforementioned approaches deal with variability of process models rather than with variability on the activities’ implementation level, on which our approach focuses.
7.2 Configurable Process Models
For managing lots of similar business processes, various approaches have emerged that allow configuring variability in process models.
The approach of hiding and blocking can be transferred to process models [52]. A reference process model contains all process variants from which concrete process models can be derived through configuration by hiding and blocking paths. A process engineer may fill in a questionnaire to configure the reference process model [32]. When an outgoing path of a process element is blocked during configuration, this path is not available any more at runtime, whereas a hidden path is skipped at runtime.
Instead of using a reference process model that contains all process variants, PROVOP [21][23][22] uses a base process model to which various change patterns may be applied in order to configure concrete process models [40]. When modelling a base process, adjustment points have to be defined to indicate where adaptions may be applied in the model. According to the PROVOP approach, possible adaptions include INSERT, DELETE and MOVE process fragments as well as MODIFY process element attributes [23]. In contrast to hiding and blocking, PROVOP allows also adding process behaviour at adjustment points during process configuration instead of just removing process elements (e.g. hiding and blocking).
Besides configuring the control flow of process models, other elements of a process model may be configurable. La Rosa et al. [31] propose an approach to individualise roles and objects modelled in a configurable process model.
For a system literature review on processes variability, the interested reader is referred to [5].
The aforementioned approaches differ from our approach as they only provide means to configure elements of the process model like paths, roles, and objects, but neglect the configuration of activity implementations, i.e. these approaches do not take an engineering perspective into account.
7.3 Software Product Line Engineering for Variable Business Processes
The concept of SPL Engineering has been applied to business processes in other works such as [50],[1], [43], [9], and [47].
In [50], the authors introduce means to mark elements in a process model as variation points. Different variants as well as their binding times may be specified as realisation for these variation points. A proof of concept is built on top of the Eclipse plugin architecture. However, the work focuses on the modelling aspect of variation points in process models rather than the implementations details of variation points. The work lacks a detailed description of the variability mechanisms used to bind the variants and a description of the management of feature interactions. In our approach, we present an in-depth approach on how to use existing variability mechanisms to bind features at compile time, start time, and runtime. Furthermore, our approach describes how to handle feature interactions when applying concepts of SPL Engineering to variable business processes.
SPL Engineering is used in [1] as well to compose business processes. The authors argue that the services called by a business process may be owned by a third-party vendor and may be highly configurable. In order to compose a business process comprising services those services are each treated as an SPL whereas the business process itself is considered a multiple SPL. The authors focus on aggregating the feature models of the SPLs, automatically checking the feature selection against the aggregated feature model, and determining which services are able to accept the output data of other services as input data. Although the authors describe how to check whether or not the services are compatible (e.g. in regard to data structures), they fail to describe how the services are automatically composed, what technical mechanism shall be used to automatically bind and connect the services. Furthermore, dynamic binding of the services and multiple binding for one activity are neglected by this work. Besides, the considered business processes only include automated activities and do not contain user tasks.
Concepts of SPL like feature modelling have been applied to cross-organisation business processes in [43]. However, the approach is not holistic and does not describe how to systematically and automatically instantiate software products derived from the SPL.
In [9], the customers’ requirements and the corresponding parts of the business process that implement these requirements are mapped in a decision table, which can be used to configure and derive a business process from the product line. However, the authors do not describe how the configuration works from a technical perspective.
SPL Engineering concepts were also applied to automatically generate process model variants. Changes in process models can be automatically propagated to the derived process model variants [47]. Our work distinguishes itself from the above one in that it focuses on the automatic generation of PAISs and not only on the automatic generation of process models.
8 Conclusions
A PAIS Product Line comprises a common set of core artefacts including a core process model. Activities of the core process model may be identified as variation points for which different implementations (e.g. user form, automated service) are developed. PAIS Products may be derived from the PAIS Product Line by selecting implementations for the variation points. Thereby, development efforts and costs are reduced when implementing similar PAIS Products sharing common features.
The presented work complemented the approach of PAIS Product Lines towards more flexibility. A plugin mechanism was introduced that is based on FeatureHouse and parametrization. Using this plugin mechanism, activity implementations may be bound at compile time, start time, and runtime. Furthermore, FeatureHouse was used to compose the process data structure. Different organisations may require varying process data structures for the same business process. By composing the process data structure using FeatureHouse, the process data structure may be adapted to an organisation’s specific needs during PAIS Product derivation. In addition, the plugin mechanism enables the selection if multiple implementations for an activity. It has been deduced that aggregation code is required when selecting multiple implementations for one activity that write to the same part of the process data structure to prevent the implementations from overwriting each others data. Aggregation code is additional code being included when multiple implementations for one activity are selected and that aggregates the output data of those implementations. Finally, FeatureIDE is used as tool support for the phases domain analysis, domain implementation, requirements analysis, and product derivation.
The presented approach has limitations, which will be addressed in future research. It focuses on varying activity implementations and assumes a fixed and immutable control flow of the process model during PAIS Product derivation. Future research shall reveal how the control flow may be adapted during PAIS Product derivation. The approaches of base and reference process models might be combined and included into the approach of PAIS Product Lines. Besides, only activities may constitute a variation point, i.e only the implementation of activities may be selected and substituted. Future research shall outline how events may constitute a variation point and how their implementation may be substituted.